OpenStudy (anonymous):

x^6-16x^4 factorize completely

jimthompson5910 (jim_thompson5910):

x^6-16x^4 (x^3)^2-(4x^2)^2 (x^3-4x^2)(x^3+4x^2) So x^6-16x^4=(x^3-4x^2)(x^3+4x^2)

OpenStudy (anonymous):

x^4(x^2-16) =x^4(x+4)(x-4)

OpenStudy (saifoo.khan):

x^4 (x-4)(x+4)